# Extended mass and spheroidal deformation effects on epicyclic frequencies and periapsis shift in quasi-circular orbits

### v1

@@ Abstract
We investigate the effects of extended mass and spheroidal deformation on the periapsis shift of quasi-circular orbits inside a gravitating mass distribution in the Newtonian framework. Focusing on the internal gravitational potential of a spheroidal body with both homogeneous and inhomogeneous density profiles, we elucidate how the ratio of local density to average density governs the extended mass effect on the periapsis shift. By analyzing the orbital angular frequency, along with the radial and vertical epicyclic frequencies, we demonstrate that in the uniform density case (i.e., the Maclaurin spheroid), where the potential takes the form of a harmonic oscillator, the periapsis exhibits a constant retrograde shift of $-\pi$. In contrast, in regions where density inhomogeneity and spheroidal deformation (in both prolate and oblate forms) are significant, the periapsis shift varies with the guiding orbital radius due to local density contrast and deformation effects. The resuls indicate that oblate deformation suppresses the extended mass effect associated with the ratio of local density to average density, whereas prolate deformation amplifies it. Furthermore, by varying the density distribution parameters, we establish the conditions for orbital stability and identify the emergence of marginally stable orbits.
@@ I Introduction
Motivated by these findings, theoretical research in this field has grown increasingly active. Various approaches have been developed to address periapsis shifts, including the post-Newtonian approximation [8, 9, 10, 11, 12], the numerical relativity [13, 14], and series expansion methods in gravitational theory [15, 16, 18, 17, 19]. Moreover, a novel expansion method has recently been introduced [20, 21]. Since quasi-circular orbits constitute the leading-order in the eccentricity expansion, a thorough understanding of them is essential for accurately determining higher-order corrections and periapsis shifts.
@@
Motivated by this background, we examine the periapsis shift in quasi-circular orbits within Newtonian theory by employing a generalized mass distribution model that extends the conventional spherically symmetric case through the incorporation of a one-parameter deformation. We investigate how this deformation influences extended mass effects by establishing a quantitative relation between the periapsis shift angle and both local and quasi-local average mass densities.
This paper is organized as follows. In Sec. II, we review extended mass effects on the periapsis shift, focusing on the local to average density ratio in a static, spherically symmetric model. In Sec. III, we derive the gravitational potential for an ellipsoidal mass distribution, with a particular focus on a solid spheroid. In Sec. IV, we formulate the motion of a test particle on quasi-circular orbits within a reflection symmetric plane, deriving the orbital angular and epicyclic frequencies, and defining the periapsis shift angle. In Sec.V, we analyze the frequency and periapsis shift properties for various density distributions and discuss the effects of source deformation. Finally, in Sec. VI, we summarize our results and outline directions for future research.
@@
In this section, we examine the periapsis shift angle of quasi-circular orbits in a static, spherically symmetric Newtonian gravitational system, and review the extended mass effects that contribute to this shift.
First, we consider a quasi-circular orbit, defined as an orbit with an negligibly small eccentricity achieved by applying a small radial perturbation to an otherwise stable circular orbit. In other words, a quasi-circular orbit oscillates harmonically in the radial direction at an epicyclic frequency $\omega_{\mathrm{r}}$, with its motion centered on the stable circular guiding orbit. Furthermore, a periapsis shift occurs when the ratio of the orbital angular frequency $\omega_{\phi}$ to the radial oscillation frequency $\omega_{\mathrm{r}}$ deviates from unity. The periapsis shift angle, which quantifies this effect, is defined as
@@ IV Epicyclic oscillations and periapsis shift in quasi-circular orbits
Here, $\rho(\xi^{2})$ and $\bar{\rho}(\xi^{2})$, defined in Eqs. (13) and (23), respectively, are both evaluated at $z=0$. Since $m^{2}=\xi^{2}$ at $z=0$, we neglect variations along the $z$-direction (e.g., due to deformation or density inhomogeneity), which results in the same form as that of the spherically symmetric case. We introduce the phenomenological parameter $f$ to account for the additional contributions to the periapsis shift arising from deformation and associated density inhomogeneities, which are absent in a purely spherically symmetric system. When $f>0$, these influences not only amplify the extended mass effect but also add a purely deformation-induced contribution that further shifts the periapsis in the negative direction, yielding a more pronounced negative shift than that produced by the density ratio alone. Conversely, when $f<0$, the deformation effects are partially canceled, thereby reducing the magnitude of the negative periapsis shift. Thus, even in regions where the local density may vanish, these effects can lead to a nonzero value of $f$. In the spherically symmetric case (i.e., $e=0$), $f$ vanishes identically, thus recovering Eq. (3).
@@ V.1 $n=0$
We begin by considering the case $n=0$, which corresponds to a spheroid with a uniform density distribution (the Maclaurin spheroid). Inside the spheroid, the particle dynamics is governed by the effective potential (26) (with $n=0$), which is given explicitly by
@@ V.2 $n=1$
This leading term is the constant $-\pi$ (independent of $e$), because the first nonconstant term in the expansion of the potential $\Phi_{1}$ in powers of $\xi$ is quadratic in $\xi$ (i.e., that of a harmonic oscillator).
@@
Figure 3 shows the $R$-dependence of $\Delta\phi_{\mathrm{p}}$ (left panel) and of the normalized parameter $f/(1+3\rho/\bar{\rho})$ (right panel) in the nonlinear regime with respect to $e^{2}$. The left panel shows that, in both the oblate (solid curves) and prolate (dashed curves) cases, $\Delta\phi_{\mathrm{p}}$ increases monotonically with $R$, consistent with the result of $\partial(\Delta\phi_{\mathrm{p}})/\partial R\geq 0$ for all defined values of $e$. In particular, for $e=0$ (i.e., $f=0$), the effect of density inhomogeneity on $\Delta\phi_{\mathrm{p}}$ evidently appears to increase from $-\pi$ toward 00, which can be interpreted as a reduction in the extended mass effect resulting from a decrease in the ratio of local to average density. For $e\neq 0$, a similar monotonic increase can be attributed to the same factors; furthermore, the additional influence of deformation arises from density inhomogeneity, an effect that is absent in the homogeneous case. For the prolate case, an increasing $e$ shifts $\Delta\phi_{\mathrm{p}}$ in the negative direction, indicating that the prolateness amplifies the extended mass effect arising from the mass density. In contrast, for the oblate case, an increasing $e$ shifts $\Delta\phi_{\mathrm{p}}$ in the positive direction, such that near the surface $\Delta\phi_{\mathrm{p}}$ becomes positive, meaning that the oblateness suppresses the extended mass effect associated with the mass density. Moreover, for $e\geq e_{0}$, as $R$ approaches $R_{\mathrm{ms}}$, $\Delta\phi_{\mathrm{p}}$ diverges to infinity because $\omega_{\mathrm{R}}$ simultaneously tends to zero. The right panel quantitatively demonstrates that the effect of deformation becomes increasingly significant as $R$ approaches the surface. These findings agree with the behavior of the periapsis shift and indicate that the effects of deformation and the resulting density inhomogeneity increase monotonically from the center to the surface of the spheroid.
@@ V.3 $n=2$
Figure 5 shows the $R$-dependence of $\Delta\phi_{\mathrm{p}}$ in the left panel and of the normalized parameter $f/(1+3\rho/\bar{\rho})$ in the right panel. In the prolate case (dashed curves), $\Delta\phi_{\mathrm{p}}$ increases monotonically with $R$ and changes in the negative direction as $e$ increases. Furthermore, the normalized $f$ indicates that the effect of deformation becomes more pronounced as $R$ increases. In contrast, in the oblate case (solid curves), $\Delta\phi_{\mathrm{p}}$ consistently changes in the positive direction as $e$ increases, but its $R$-dependence varies qualitatively with $e$. The normalized $f$ further suggests that the deformation effect suppresses the extended mass effect.
The behavior of $\Delta\phi_{\mathrm{p}}$ undergoes a significant transition at the critical value $e_{0}=0.988937349$. The left panel of Fig. 6 shows the region in the $R$-$e$ plane where quasi-circular orbits exist (blue-shaded) and where they do not (unshaded). When $e<e_{0}$, quasi-circular orbits exist over the entire range of $R$, whereas for $e_{0}\leq e<1$, a region in $R$ emerges where quasi-circular orbits are absent (i.e., where $\omega^{2}_{\mathrm{R}}<0$). Note that, unlike in the $n=1$ case where the marginally stable circular orbit appears at the spheroid’s surface, at $e=e_{0}$ it occurs at $R/a=0.969442391$. Consequently, for $e$ in the range $[e_{0},1)$, $\Delta\phi_{\mathrm{p}}$ diverges at those radii corresponding to the two marginally stable circular orbits (see the right panel of Fig. 6).
In the oblate case, we focus on $e=0.9$ as shown in Fig. 5. Here, $\Delta\phi_{\mathrm{p}}$ exhibits a peak and then decreases in the surface region. This behavior is partly because the extended mass effect weakens in the surface region due to the low local density. Similarly, the normalized function $f/(1+3\rho/\bar{\rho})$ with negative values also exhibits a minimum near the surface before increasing with $R$, which implies a reduced contribution from the deformation effect. The combined effects of these factors determine the behavior of the shift angle in the near-surface region.
@@ VI Summary and discussion
Within the framework of Newtonian theory, we have investigated the effects of both extended mass and gravitational source deformation on the periapsis shift phenomenon in quasi-circular orbits. Specifically, we focused on an ellipsoidal source consisting of uniform-density homeoid layers. By assuming a specific density profile and a spheroidal configuration, we analyzed the quasi-circular orbits confined to the reflection-symmetric plane of the interior gravitational potential. Using this approach, we derived expressions for the orbital angular frequency, the epicyclic frequencies, and the periapsis shift angle. Furthermore, by varying the parameters inherent in the density distribution, we provided a detailed analysis of how spheroidal deformation and associated density inhomogeneities affect these frequencies and the shift angle.
@@
We show that the effects of extended mass and source deformation on quasi-circular orbits vary significantly with the degree of density inhomogeneity and the extent of spheroidal deformation. In the case of uniform density, the internal gravitational potential reduces to that of a harmonic oscillator; consequently, both the orbital angular frequency and epicyclic frequencies become independent of the guiding orbit radius, and the periapsis shift angle remains constant at $-\pi$. This implies that for a uniform-density spheroid, the deformation does not affect the periapsis shift. In contrast, for inhomogeneous density distributions, both the orbital angular frequency and the epicyclic frequencies depend on the guiding orbit radius and the deformation parameter, causing the shift angle to deviate from the uniform-density value. As in the spherically symmetric system, the extended mass effect influences the shift angle through the ratio of the local density to the average density, with a higher local density contributing to a more negative shift. Moreover, we find that in the case of oblate deformation, the extended mass effect associated with the ratio of local to average mass density is suppressed, whereas in the case of prolate deformation it is amplified.
@@
We also analyzed the stability of quasi-circular orbits in this system by examining the epicyclic frequencies. In the uniform-density case, the quasi-circular orbits remain stable regardless of the guiding orbit radius or the deformation parameter. In contrast, when the density is inhomogeneous—especially under oblate deformation—the radial epicyclic frequency becomes zero, indicating that marginally stable circular orbits appear for $e_{0}\leq e<1$. This phenomenon seems to occur only for $n=1$ and $n=2$. Specifically, for $n=1$, the critical value is $e_{0}=0.929955685$, which agrees with the oblateness at which marginally stable circular orbits emerge in the external vacuum region [47]. This agreement occurs because, in both the interior region and the external vacuum region, the marginally stable circular orbit emerges at the spheroid’s surface when $e=e_{0}$. In contrast, for $n=2$, the critical value is $e_{0}=0.988937349$, and no such correspondence with the external vacuum region’s oblateness is observed. Notably, the marginally stable orbits—often misinterpreted as a phenomenon unique to general relativity—actually emerge under Newtonian conditions due to extended mass effects and density inhomogeneity. Meanwhile, under prolate deformation, marginally stable orbits are not formed.
@@
We will report the results of our post-Newtonian analysis in a separate paper, which examines the competition between the general-relativistic effects and the combined extended mass and deformation effects on the periapsis shift of quasi-circular orbits in a system featuring a black hole at the center of a spheroidal mass distribution. Furthermore, the extended mass and deformation effects on the periapsis shift of quasi-circular orbits in static, axisymmetric spacetimes warrant future work. In addition, generalizing the general formula for the periapsis shift of quasi-circular orbits in static, spherically symmetric systems (2) to static, axisymmetric spacetimes is a promising topic for future research.

### v2

@@ Abstract
We investigate the effects of extended mass and spheroidal deformation on the periapsis shift of quasi-circular orbits inside a gravitating mass distribution in the Newtonian framework. Focusing on the internal gravitational potential of a spheroidal body with both homogeneous and inhomogeneous density profiles, we elucidate how the ratio of local density to average density governs the extended mass effect on the periapsis shift. By analyzing the orbital angular frequency, along with the radial and vertical epicyclic frequencies, we demonstrate that in the uniform density case (i.e., the Maclaurin spheroid), where the potential takes the form of a harmonic oscillator, the periapsis exhibits a constant retrograde shift of $-\pi$. In contrast, in regions where density inhomogeneity and spheroidal deformation (in both prolate and oblate forms) are significant, the periapsis shift varies with the guiding orbital radius due to local density contrast and deformation effects. The results indicate that oblate deformation suppresses the extended mass effect associated with the ratio of local density to average density, whereas prolate deformation amplifies it. Furthermore, by varying the density distribution parameters, we establish the conditions for orbital stability and identify the emergence of marginally stable orbits.
@@ I Introduction
Motivated by these findings, theoretical research in this field has grown increasingly active. Various approaches have been developed to address periapsis shifts, including the post-Newtonian approximation [8, 9, 10, 11, 12], the numerical relativity [13, 14], and series expansion methods in gravitational theory [15, 16, 18, 17, 19]. Moreover, a novel expansion method has recently been introduced [20, 21]. Since quasi-circular orbits constitute the leading order in the eccentricity expansion, a thorough understanding of them is essential for accurately determining higher-order corrections and periapsis shifts.
@@
Motivated by this background, we examine the periapsis shift in quasi-circular orbits within Newtonian theory by employing a generalized mass distribution model that extends the conventional spherically symmetric case by incorporating a one-parameter deformation. We investigate how this deformation influences extended mass effects by establishing a quantitative relation between the periapsis shift angle and both local and quasi-local average mass densities.
This paper is organized as follows. In Sec. II, we review extended mass effects on the periapsis shift, focusing on the local-to-average density ratio in a static, spherically symmetric model. In Sec. III, we derive the gravitational potential for an ellipsoidal mass distribution, focusing on a solid spheroid. In Sec. IV, we formulate the motion of a test particle on quasi-circular orbits within a reflection symmetric plane, deriving the orbital angular and epicyclic frequencies, and defining the periapsis shift angle. In Sec.V, we analyze the frequency and periapsis shift properties for various density distributions and discuss the effects of source deformation. Finally, in Sec. VI, we summarize our results and outline directions for future research.
@@
In this section, we examine the periapsis shift angle of quasi-circular orbits in a static, spherically symmetric Newtonian gravitational system, and review the extended mass effects contributing to this shift.
First, we consider a quasi-circular orbit, defined as an orbit with a negligibly small eccentricity achieved by applying a small radial perturbation to an otherwise stable circular orbit. In other words, a quasi-circular orbit oscillates harmonically in the radial direction at an epicyclic frequency $\omega_{\mathrm{r}}$, with its motion centered on the stable circular guiding orbit. Furthermore, a periapsis shift occurs when the ratio of the orbital angular frequency $\omega_{\phi}$ to the radial oscillation frequency $\omega_{\mathrm{r}}$ deviates from unity. The periapsis shift angle, which quantifies this effect, is defined as
@@ IV Epicyclic oscillations and periapsis shift in quasi-circular orbits
Here, $\rho(\xi^{2})$ and $\bar{\rho}(\xi^{2})$, defined in Eqs. (13) and (23), respectively, are both evaluated at $z=0$. Since $m^{2}=\xi^{2}$ at $z=0$, we neglect variations along the $z$-direction (e.g., due to deformation or density inhomogeneity), resulting in the same form as the spherically symmetric case. We introduce the phenomenological parameter $f$ to account for the additional contributions to the periapsis shift arising from deformation and associated density inhomogeneities, which are absent in a purely spherically symmetric system. When $f>0$, these influences not only amplify the extended mass effect but also add a purely deformation-induced contribution that further shifts the periapsis in the negative direction, yielding a more pronounced negative shift than that produced by the density ratio alone. Conversely, when $f<0$, the deformation effects are partially canceled, thereby reducing the magnitude of the negative periapsis shift. Thus, even in regions where the local density may vanish, these effects can lead to a nonzero value of $f$. In the spherically symmetric case (i.e., $e=0$), $f$ vanishes identically, thus recovering Eq. (3).
@@ V.1 $n=0$
We begin by considering the case $n=0$, corresponding to a spheroid with a uniform density distribution (the Maclaurin spheroid). Inside the spheroid, the particle dynamics is governed by the effective potential (26) (with $n=0$), which is given explicitly by
@@ V.2 $n=1$
This leading term is the constant $-\pi$ (independent of $e$) because the first nonconstant term in the expansion of the potential $\Phi_{1}$ in powers of $\xi$ is quadratic in $\xi$ (i.e., that of a harmonic oscillator).
@@
Figure 3 shows the $R$-dependence of $\Delta\phi_{\mathrm{p}}$ (left panel) and of the normalized parameter $f/(1+3\rho/\bar{\rho})$ (right panel) in the nonlinear regime with respect to $e^{2}$. The left panel shows that, in both the oblate (solid curves) and prolate (dashed curves) cases, $\Delta\phi_{\mathrm{p}}$ increases monotonically with $R$, consistent with the result of $\partial(\Delta\phi_{\mathrm{p}})/\partial R\geq 0$ for all defined values of $e$. In particular, for $e=0$ (i.e., $f=0$), the effect of density inhomogeneity on $\Delta\phi_{\mathrm{p}}$ evidently appears to increase from $-\pi$ toward 00, which can be interpreted as a reduction in the extended mass effect resulting from a decrease in the ratio of local to average density. For $e\neq 0$, a similar monotonic increase can be attributed to the same factors; furthermore, the additional influence of deformation arises from density inhomogeneity, an effect that is absent in the homogeneous case. For the prolate case, an increasing $e$ shifts $\Delta\phi_{\mathrm{p}}$ in the negative direction, indicating that the prolateness amplifies the extended mass effect arising from the mass density. In contrast, for the oblate case, an increasing $e$ shifts $\Delta\phi_{\mathrm{p}}$ in the positive direction, such that near the surface $\Delta\phi_{\mathrm{p}}$ becomes positive, meaning that the oblateness suppresses the extended mass effect associated with the mass density. Moreover, for $e\geq e_{0}$, as $R$ approaches $R_{\mathrm{ms}}$, $\Delta\phi_{\mathrm{p}}$ diverges to infinity because $\omega_{\mathrm{R}}$ simultaneously tends to zero. The right panel quantitatively demonstrates that the deformation effect becomes increasingly significant as $R$ approaches the surface. These findings agree with the behavior of the periapsis shift and indicate that the effects of deformation and the resulting density inhomogeneity increase monotonically from the center to the surface of the spheroid.
@@ V.3 $n=2$
Figure 5 shows the $R$-dependence of $\Delta\phi_{\mathrm{p}}$ in the left panel and of the normalized parameter $f/(1+3\rho/\bar{\rho})$ in the right panel. In the prolate case (dashed curves), $\Delta\phi_{\mathrm{p}}$ increases monotonically with $R$ and changes in the negative direction as $e$ increases. Furthermore, the normalized parameter indicates that the deformation effect becomes more pronounced as $R$ increases. In contrast, in the oblate case (solid curves), $\Delta\phi_{\mathrm{p}}$ consistently changes in the positive direction as $e$ increases, but its $R$-dependence varies qualitatively with $e$. The normalized parameter further suggests that the deformation effect suppresses the extended mass effect.
The behavior of $\Delta\phi_{\mathrm{p}}$ undergoes a significant transition at the critical value $e_{0}=0.988937349$. The left panel of Fig. 6 shows the region in the $R$-$e$ plane where quasi-circular orbits exist (blue-shaded) and where they do not (unshaded). When $e<e_{0}$, quasi-circular orbits exist over the entire range of $R$, whereas for $e_{0}\leq e<1$, a region in $R$ emerges where quasi-circular orbits are absent (i.e., where $\omega^{2}_{\mathrm{R}}<0$). Note that, unlike in the $n=1$ case where the marginally stable circular orbit appears at the spheroid’s surface, at $e=e_{0}$, it occurs at $R/a=0.969442391$. Consequently, for $e$ in the range $[e_{0},1)$, $\Delta\phi_{\mathrm{p}}$ diverges at those radii corresponding to the two marginally stable circular orbits (see the right panel of Fig. 6).
In the oblate case, we focus on $e=0.9$ as shown in Fig. 5. Here, $\Delta\phi_{\mathrm{p}}$ exhibits a peak and then decreases near the surface. This behavior is partly because the extended mass effect weakens in the surface region due to the low local density. Similarly, the normalized function $f/(1+3\rho/\bar{\rho})$ with negative values also exhibits a minimum near the surface before increasing with $R$, indicating a reduced contribution from the deformation effect. The combined effects of these factors determine the behavior of the shift angle in the near-surface region.
@@ VI Summary and discussion
Within the framework of Newtonian theory, we have investigated the effects of both extended mass and gravitational source deformation on the periapsis shift phenomenon in quasi-circular orbits. Specifically, we focused on an ellipsoidal source consisting of uniform-density homeoid layers. By assuming a specific density profile and a spheroidal configuration, we analyzed the quasi-circular orbits confined to the reflection-symmetric plane of the interior gravitational potential. Using this approach, we derived expressions for the orbital angular frequency, the epicyclic frequencies, and the periapsis shift angle. Furthermore, by varying the parameters inherent in the density distribution, we provide a detailed analysis of how spheroidal deformation and associated density inhomogeneities affect these frequencies and the shift angle.
@@
We show that the effects of extended mass and source deformation on quasi-circular orbits vary significantly with the degree of density inhomogeneity and the extent of spheroidal deformation. In the case of uniform density, the internal gravitational potential reduces to that of a harmonic oscillator; consequently, both the orbital angular frequency and epicyclic frequencies become independent of the guiding orbit radius, and the periapsis shift angle remains constant at $-\pi$. This implies that for a uniform-density spheroid, the deformation does not affect the periapsis shift. In contrast, for inhomogeneous density distributions, both the orbital angular frequency and the epicyclic frequencies depend on the guiding orbit radius and the deformation parameter, causing the shift angle to deviate from the uniform-density value. As in the spherically symmetric system, the extended mass effect influences the shift angle through the ratio of the local density to the average density, with a higher local density contributing to a more negative shift. Moreover, we find that in the case of oblate deformation, this effect is suppressed, whereas in the case of prolate deformation, it is amplified.
@@
We also analyzed the stability of quasi-circular orbits in this system by examining the epicyclic frequencies. In the uniform-density case, the quasi-circular orbits remain stable regardless of the guiding orbit radius or the deformation parameter. In contrast, when the density is inhomogeneous—especially under oblate deformation—the radial epicyclic frequency becomes zero, indicating that marginally stable circular orbits appear for $e_{0}\leq e<1$. This phenomenon seems to occur only for $n=1$ and $n=2$. Specifically, for $n=1$, the critical value is $e_{0}=0.929955685$, which agrees with the oblateness at which marginally stable circular orbits emerge in the external vacuum region [47]. This agreement occurs because, in both the interior region and the external vacuum region, the marginally stable circular orbit emerges at the spheroid’s surface when $e=e_{0}$. In contrast, for $n=2$, the critical value is $e_{0}=0.988937349$, and no such correspondence with the external vacuum region’s oblateness is observed. Notably, the marginally stable orbits—often misinterpreted as a phenomenon unique to general relativity—emerge under Newtonian conditions due to extended mass effects and density inhomogeneity. Meanwhile, under prolate deformation, marginally stable orbits are not formed.
@@
We will report the results of our post-Newtonian analysis in a separate paper, which examines the competition between the general-relativistic effects and the combined extended mass and deformation effects on the periapsis shift of quasi-circular orbits in a system featuring a black hole at the center of a spheroidal mass distribution. Furthermore, the extended mass and deformation effects on the periapsis shift of quasi-circular orbits in static, axisymmetric spacetimes warrant future work [51]. In addition, generalizing the general formula for the periapsis shift of quasi-circular orbits in static, spherically symmetric systems (2) to static, axisymmetric spacetimes is a promising topic for future research.